# Gravitational lensing inside and outside of a marginally unstable photon sphere in a general, static, spherically symmetric, and asymptotically-flat spacetime in strong deflection limits

### v1

@@ IV Observable of gravitational lensing in the strong deflection limits
Every gravitational lensed image with the winding number $n\geq 1$ and with a positive impact parameter makes a pair of the image with the same winding number $n\geq 1$ and with a negative impact parameter. The pair of images with every winding number $n\geq 1$ has the almost same absolute values of the image angles and the magnifications but a different signs under the assumptions of the small angles. Thus, the diameter of the paired images is given by $2\theta_{n\pm}$ with the winding number $n\geq 1$ for the positive impact parameter and the total magnification $\mu_{n\mathrm{tot}\pm}$, which is defined by sum of the absolute value of the magnification of the pair of the images with the winding number $n\geq 1$, is given by
@@
In Table I, the observable with the distances $d_{\mathrm{OS}}=16$kpc and $d_{\mathrm{OL}}=d_{\mathrm{LS}}=8$kpc, the mass $M=4\times 10^{6}M_{\odot}$ in the Reissner-Nordström spacetime and the Hayward spacetime is shown for the winding numbers $n=1$, 22, and $3$.
@@
$\mu_{1\mathrm{tot}+}(\phi)\times 10^{19}$ 1063.82 1063.19 324.595 890.522 - 271.347
$\mu_{2\mathrm{tot}+}(\phi)\times 10^{19}$ 52.8675 52.8477 16.4542 44.996 - 13.9129
$\mu_{3\mathrm{tot}+}(\phi)\times 10^{19}$ 6.58948 6.58765 2.05397 5.62252 - 1.7433
$\bar{r}_{+}$ 17.3389 17.3350 16.9947 17.0469 - 16.7917
$\bar{c}_{-}$ 11.5658 11.5657 - 10.4155 - -
$\bar{d}_{-}$ -5.64071 -5.59108 - -5.69675 - -
@@
$\mu_{1\mathrm{tot}-}(\phi)\times 10^{19}$ 4518.59 4066.21 - 12739.8 - -
$\mu_{2\mathrm{tot}-}(\phi)\times 10^{19}$ 1304.72 1328.28 - 1131.53 - -
$\mu_{3\mathrm{tot}-}(\phi)\times 10^{19}$ 173.347 175.793 - 147.045 - -
$\bar{r}_{-}$ 2.95597 2.61464 - 9.64064 - -
V Discussion and Conclusion
If the relation $\bar{d}_{-}=\bar{d}_{+}$ is also correct in other spacetimes, the relation would be used to check our numerical method. In the Hayward spacetime with the marginally unstable photon sphere, we have obtained $\bar{d}_{+}\sim-5.60584$ and $\bar{d}_{-}\sim-5.69675$ in our method and they imply that the relation $\bar{d}_{-}=\bar{d}_{+}$ is almost held. We have approximately read $r_{0}\rightarrow r_{\mathrm{m}}+0$ in Eqs. (28) and (29) as $r_{0}=r_{\mathrm{m}}+5\times 10^{-5}M$ for numerical calculations to get the values of $\bar{c}_{+}$ and $\bar{d}_{+}$ and we have used $r_{0}=r_{\mathrm{m}}-5\times 10^{-5}M$ for numerical calculations of $r_{0}\rightarrow r_{\mathrm{m}}-0$ in Eqs. (28) and (29) to get the values of $\bar{c}_{-}$ and $\bar{d}_{-}$.
In the Hayward spacetime with the marginally unstable photon sphere, we have obtained $\bar{c}_{+}\sim 6.01322$ in our numerical method and its difference from the value of $\bar{c}_{+}\sim 6.01316$ in Chiba and Kimura Chiba:2017nml is tiny but we have found that the value $\bar{c}_{+}\sim 4.95196$ obtained by the semi-analytic calculation in Tsukamoto:2020iez should be modified.
Thus, we conclude that our extent Eiroa, Romero, and Torres’s method to gravitational lensing of the rays just inside and outside of a marginally unstable photon sphere in a general, static, spherically symmetric, and asymptotically-flat spacetime in strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$ gives the correct values of $\bar{c}_{\pm}$ and $\bar{d}_{\pm}$ and that the semi-analytic method considered by the author previously in Ref. Tsukamoto:2020iez gives the invalid values of $\bar{c}_{+}$ while it gives the correct values of $\bar{d}_{+}$.
We comment on constraints on metrics from the shadow observations of supermassive black hole candidates at the centers of our galaxy and the elliptic galaxy M87 by the Event Horizon Telescope Akiyama:2019cqa ; EventHorizonTelescope:2022wkp . The Event Horizon Telescope collaborations EventHorizonTelescope:2021dqv ; EventHorizonTelescope:2022xqj constrained the parameter of the metrics of spacetimes from the observations and the sizes of the photon spheres under assumptions for simplify. Thus, the constraints. which depend on the assumptions. are not very reliable Gralla:2020pra ; Tsukamoto:2024gkz and the constraints on the metrics should be confirmed by near-future space observations Lupsasca:2024xhq ; Johnson:2024ttr .

### v2

@@ IV Observable of gravitational lensing in the strong deflection limits
Every gravitational lensed image with the winding number $n\geq 1$ and with a positive impact parameter makes a pair of the image with the same winding number $n\geq 1$ and with a negative impact parameter. The pair of images with every winding number $n\geq 1$ has the almost same absolute values of the image angles and the magnifications but a different signs under the assumptions of the small angles. Thus, the diameter of the paired images is given by $2\theta_{n\pm}$ with the winding number $n\geq 1$ for the positive impact parameter and the total magnification $\mu_{n\mathrm{tot}\pm}$, which is defined by the sum of the absolute values of the magnification of the pair of the images with the winding number $n\geq 1$, is given by
@@
V Conclusion
If the relation $\bar{d}_{-}=\bar{d}_{+}$ is also correct in other spacetimes, the relation would be used to check our numerical method. In the Hayward spacetime with the marginally unstable photon sphere, we have obtained $\bar{d}_{+}\sim-5.60958$ and $\bar{d}_{-}\sim-5.67374$ in our method and they imply that the relation $\bar{d}_{-}=\bar{d}_{+}$ is almost held. We have approximately read $r_{0}\rightarrow r_{\mathrm{m}}+0$ in Eqs. (29) and (30) as $r_{0}=r_{\mathrm{m}}+3\times 10^{-5}M$ for numerical calculations to get the values of $\bar{c}_{+}$ and $\bar{d}_{+}$ and we have used $r_{0}=r_{\mathrm{m}}-2.4\times 10^{-5}M$ for numerical calculations of $r_{0}\rightarrow r_{\mathrm{m}}-0$ in Eqs. (29) and (30) to get the values of $\bar{c}_{-}$ and $\bar{d}_{-}$.
In the Hayward spacetime with the marginally unstable photon sphere, we have obtained $\bar{c}_{+}\sim 6.01324$ in our numerical method and its difference from the value of $\bar{c}_{+}\sim 6.01316$ in Chiba and Kimura Chiba:2017nml is tiny but we have found that the value $\bar{c}_{+}\sim 4.95196$ obtained by the semi-analytic calculation in Ref. Tsukamoto:2020iez should be modified.
Thus, we conclude that our extent Eiroa, Romero, and Torres’s method to gravitational lensing of the rays just inside and outside of a marginally unstable photon sphere in a general, static, spherically symmetric, and asymptotically-flat spacetime in strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$ gives the correct values of $\bar{c}_{\pm}$ and $\bar{d}_{\pm}$ and that the semi-analytic method considered by the author previously in Ref. Tsukamoto:2020iez gives the invalid values of $\bar{c}_{+}$ while it gives the correct values of $\bar{d}_{+}$. We will revisit to modify the semi-analytic calculation to obtain $\bar{c}_{\pm}$ correctly in forthcoming work Igata:2026ivq .
On this paper, we have concentrated on the extension of Eiroa, Romero, and Torres’s method to gravitational lensing of the rays inside and outside of the marginally unstable photon sphere. One may expect that we can apply it to the deflection angles (2) in the strong deflection limits $b\rightarrow b_{\mathrm{m}}\pm 0$ in spacetimes with a photon sphere and an antiphoton sphere. As shown in Refs. Shaikh:2019itn ; Tsukamoto:2021fsz , however, $r_{0}$ does not approach to $r_{\mathrm{m}}$ but another value in the limit $b\rightarrow b_{\mathrm{m}}-0$. Thus, the strong deflection limit $r_{0}\rightarrow r_{\mathrm{m}}-0$ in Eqs. (27)-(30) is invalid and we need another extension of Eiroa, Romero, and Torres’s method for the deflection angles (2) in the strong deflection limit $b\rightarrow b_{\mathrm{m}}-0$ and it is left as a future work.
VI Discussion
The observed ring images of the supermassive black hole candidates at the centers of our galaxy and the elliptic galaxy M87 by the Event Horizon Telescope Akiyama:2019cqa ; EventHorizonTelescope:2022wkp are consistent with lensing images formed by synchrotron radiations from a hot plasma near the black holes according to their ray-traced general-relativistic magnetohydrodynamic (GRMHD) simulations but the exotic compact objects such as naked singularity spacetimes, regular compact objects, and wormholes are not excluded Akiyama:2019cqa ; EventHorizonTelescope:2022wkp . This is because there are too many parameter regions for the exotic compact objects Akiyama:2019cqa ; EventHorizonTelescope:2022wkp and astrophisical models Gralla:2020pra to exclude the exotic cases.
We comment on constraints on metrics from the shadow observations of supermassive black hole candidates at the centers of our galaxy and M87 by the Event Horizon Telescope Akiyama:2019cqa ; EventHorizonTelescope:2022wkp . The Event Horizon Telescope collaborations EventHorizonTelescope:2021dqv ; EventHorizonTelescope:2022xqj constrained the parameter of the metrics of spacetimes from the observations and the sizes of the photon spheres under assumptions for simplify. Thus, the constraints, which depend on the assumptions are not very reliable Gralla:2020pra ; Tsukamoto:2024gkz and the constraints on the metrics should be confirmed by near-future space observations Lupsasca:2024xhq ; Johnson:2024ttr .
[…]
A compact object with a marginary unstable photon sphere can be interesting exotic cases since it can be one of the most different cases from a Schwarzschild black hole among compact objects with diverging deflection angles. For the comparisons of the observable, we have to treat the rays inside of the marginally unstable photon sphere also since they can brighter than rays nearly outside the marginally unstable photon sphere. In Refs. Tsukamoto:2020iez and Chiba:2017nml , they do not give formulas for inside of the marginary unstable photon sphere while we and Ref. Sasaki:2025web do that. In Ref. Sasaki:2025web , we can see a complete and reliable treatment in the Reissner-Nordström spacetime case. Our approach in this paper will be gives larger errors than Ref. Sasaki:2025web but it has a merit that it is easy to apply to other spacetimes. The precise values in Ref Sasaki:2025web in the Reissner-Nordström spacetime case will be a milestone to improve our method more. We, however, emphasize that the larger errors in our numerical values in Table I does not affect our discussion as we seen below.
[…]
@@ VI Discussion
$\mu_{1\mathrm{tot}+}(\phi)\times 10^{19}$ 1063.82 1063.19 324.595 888.554 - 271.347
$\mu_{2\mathrm{tot}+}(\phi)\times 10^{19}$ 52.8675 52.8477 16.4542 44.9320 - 13.9129
$\mu_{3\mathrm{tot}+}(\phi)\times 10^{19}$ 6.58948 6.58765 2.05397 5.61662 - 1.7433
$\bar{r}_{+}$ 17.3389 17.3350 16.9947 17.0324 - 16.7917
$\bar{c}_{-}$ 11.5658 11.5657 - 10.4154 - -
$\bar{d}_{-}$ -5.64071 -5.59108 - -5.67374 - -
@@
$\mu_{1\mathrm{tot}-}(\phi)\times 10^{19}$ 4518.59 4066.21 - 12798.4 - -
$\mu_{2\mathrm{tot}-}(\phi)\times 10^{19}$ 1304.72 1328.28 - 1141.19 - -
$\mu_{3\mathrm{tot}-}(\phi)\times 10^{19}$ 173.347 175.793 - 148.000 - -
$\bar{r}_{-}$ 2.95597 2.61464 - 9.60632 - -
Acknowledgements
The author would like to be grateful to T. Igata, T. Sasaki, H. Asada, and an anonymous referee for their useful and valuable comments. The author also would like to thank to M. Saijo for his comments in the very early stage of this work.